Use the following statistics to answer questions 8-13 about population parameters. Use a 5% significance level and a two tail test for all problems and specify the statistical hypotheses. Gender-0 for males; gender-1 for females correlati coefficen-100 republic demonc indepen 100 $63 $57 $49 $14 $11 $12 rats dents nt earning ans on sample size ave eam 100 100 party s party eaning 0.1 sd MMGW 7510. [3] Are the average earnings of all Republicans equal to $60?
Answer #1

Question 10

Here sample size =n = 100

average earning = $ 63

standard deviation of earning = $ 14

standard error of earning = 14/sqrt(100) = $ 1.4

Hypothesis :

H0 ; \mu = $ 60

Ha : \mu\neq $ 60

Test statistic

t = (63 - 60)/1.4 = 2.143

Here p - value = 2 * Pr(t > 2.143; dF = 99) = 0.0346 < 0.05

so we reject the nulll hypothesis and conclude that average earning of all republicans is not equal to $ 60.